# Lung-DDPM: Semantic Layout-guided Diffusion Models for Thoracic CT Image Synthesis

### v1

@@ 6 Conclusion and future study
In this paper, we introduced Lung-DDPM, a DDPM-based thoracic CT image synthesis method capable of generating 3D synthetic CT images with high diversity and fidelity. The proposed method incorporates two key components for conditional thoracic CT image synthesis: a novel anatomically aware sampling (AAS) process and an adjusted 3D U-Net denoiser structure. Through the proposed AAS process, Lung-DDPM can generate realistic thoracic CT images with precise anatomical structures and acceptable diversity. Additionally, Lung-DDPM effectively addresses three prevalent issues in SOTA generative models when performing the conditional thoracic CT image synthesis task: anatomical imprecision, imperfect blending, and inconsistency. The experimental results indicate that the proposed method outperforms other SOTA generative models in both image quality and downstream lung nodule segmentation evaluations. This highlights its strong capability to generate high-quality synthetic CT images that can be beneficial in various downstream diagnostic tasks. In our future research, we plan to redesign Lung-DDPM’s architecture for high-resolution and flexible CT image synthesis to further improve synthetic data quality.
@@ 7 Acknowledgments
Venkata Manem holds a salary support award from the IVADO and was supported by the New Frontier Research - Rapid Response Fund. This work is funded by the IUCPQ Biobank and SynergiQC.

### v2

@@ 6 Conclusion and future study
In this paper, we introduced Lung-DDPM, a DDPM-based thoracic CT image synthesis method capable of generating 3D synthetic CT images with high diversity and fidelity. The proposed method incorporates two key components for conditional thoracic CT image synthesis: a novel anatomically aware sampling (AAS) process and an adjusted 3D U-Net denoiser structure. Through the proposed AAS process, Lung-DDPM can generate realistic thoracic CT images with precise anatomical structures and acceptable diversity. Additionally, Lung-DDPM effectively addresses three prevalent issues in SOTA generative models when performing the conditional thoracic CT image synthesis task: anatomical imprecision, imperfect blending, and inconsistency. The experimental results indicate that the proposed method outperforms other SOTA generative models in both image quality and downstream lung nodule segmentation evaluations. This highlights its strong capability to generate high-quality synthetic CT images that can be beneficial in various downstream diagnostic tasks. Although Lung-DDPM provides high-quality volumetric synthesis and spatial consistency, we recognize that computational efficiency and scalability are crucial for practical deployment. In future work, we plan to conduct a comprehensive study on the trade-offs between synthesis quality and computational demands, and to further optimize the Lung-DDPM framework for memory and runtime efficiency.
@@ 7 Acknowledgments
Venkata Manem holds a salary support award from the IVADO and was supported by the New Frontier Research - Rapid Response Fund. This work is funded by the IUCPQ Biobank and foundation, and the SynergiQC project/CQDM.